# Spiking mode-based neural networks

### v2

@@ IV Conclusion
In this work, we propose a spiking mode-based neural network framework for training various computational tasks. This framework is based on mode decomposition learning inspired from the Hopfield network and multi-layered perceptron training [24, 26]. From the SMNN learning rule, we can adapt the mode size to the task difficulty, and retrieve the power-law behavior of the importance scores, and furthermore, the high dimensional recurrent activity can be projected to the low-dimensional mode space with a few leading modes, derived from the pow-law behavior. Using a few modes, we can speed up the training of recurrent spiking networks, thereby making a large scale of spike-based computation possible in practice. Further extension of our work allows us to treat more biological networks, e.g., considering excitatory-inhibitory networks, sparsity of network connections, and sequence memory from network activity, etc. It would be very interesting to see whether a biological plausible rule can be derived for a sparse sign-constrained spiking network, which we leave for future works.
To conclude, our work provides a fast, interpretable and biologically plausible framework for analyzing neuroscience experiments and designing brain-like computation circuits, and in particular, the derived plasticity rule in high dimensional spiking dynamics is linked to intriguing physics pictures of attractors, power-law behavior, which would help to elucidate inner workings of high dimensional brain dynamics, a central topic of non-equilibrium physics.

### v3

@@ IV Conclusion
In this work, we propose a spiking mode-based neural network framework for training various computational tasks. This framework is based on mode decomposition learning inspired from the Hopfield network and multi-layered perceptron training [24, 26]. From the SMNN learning rule, we can adapt the mode size to the task difficulty, and retrieve the power-law behavior of the importance scores, and furthermore, the high dimensional recurrent activity can be projected to the low-dimensional mode space with a few leading modes, derived from the power-law behavior. Using a few modes, we can speed up the training of recurrent spiking networks, thereby making a large scale of spike-based computation possible in practice. Further extension of our work allows us to treat more biological networks, e.g., considering excitatory-inhibitory networks, sparsity of network connections, and sequence memory from network activity, etc. It would be very interesting to see whether a biological plausible rule can be derived for a sparse sign-constrained spiking network, which we leave for future works.
To conclude, our work provides a fast, interpretable and biologically plausible framework for analyzing neuroscience experiments and designing brain-like computation circuits, and in particular, the derived plasticity rule in high dimensional spiking dynamics is linked to intriguing physics pictures of attractors, power-law behavior, which would help to elucidate inner workings of high dimensional brain dynamics—a central topic of non-equilibrium physics.